OpenStudy (kanwal32):

If x^2y^3=6 then find the least value of 3x+4y

OpenStudy (anonymous):

oh no not again!!

OpenStudy (anonymous):

just saw this yesterday, maybe the day before AM GM comparison

OpenStudy (anonymous):

the gimmick is to write \[3x=\frac{3x}{2}+\frac{3x}{2}\] and \[4y=\frac{4y}{3}+\frac{4y}{3}+\frac{4y}{3}\]

OpenStudy (anonymous):

the arithmetic mean of those numbers is \(\frac{3x+4y}{5}\) i will let you compute the geometric mean

OpenStudy (kanwal32):

i did the same way but was not getting the answer @ganeshie8

OpenStudy (kanwal32):

@satellite73

OpenStudy (kanwal32):

i did the same way

OpenStudy (anonymous):

that arithmetic mean of those 5 numbers is \[\frac{3x+4y}{5}\] right?

OpenStudy (kanwal32):

yes

OpenStudy (kanwal32):

i similarly broke it

OpenStudy (anonymous):

the geometric mean is ?

OpenStudy (kanwal32):

can we use cauchy shwarze property here

OpenStudy (anonymous):

oh not that complicated compute the geometric mean

OpenStudy (kanwal32):

root6

OpenStudy (anonymous):

i think it is 2

OpenStudy (kanwal32):

how????????????

OpenStudy (anonymous):

\[\sqrt[5]{\frac{3x}{2}\times \frac{3x}{2}\times \frac{4y}{3}\times \frac{4y}{3}\times\frac{4y}{3}}\]

OpenStudy (kanwal32):

ok

OpenStudy (kanwal32):

thnx @satellite73

OpenStudy (anonymous):

yw

OpenStudy (kanwal32):

ok